A series RLC circuit has a capacitor with a capacitance of 31.0 μF , an inductor with an inductance of 0.600 H and a resistor with a resistance of 88.0 Ω. The circuit has a rms current of 9.00 A when the frequency is 71.0 Hz. What is εrms?
Answer #1

Xl = 2 pi x 71 x 0.6 = 267.66 ohm

Xc = 1/2 pi x 71 x 31 x 10^-6 = 72.31 ohm

Impedance, Z = sqrt (88^2 + (267.66 - 72.31)^2) = 214.26 ohm

Emf, E = 9 x 214.26 = 1928.30 V

Comment in case any doubt please rate my answer...